SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Jun 21, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Yaw Ming Chia, Al Eisner, Fang Fang, Selina Li, Dan Walker.

Operations

  • We haven't been taking physics data since last Saturday owl shift due to PEP vacuum leak in LER. Estimation for resuming data-taking is this Thursday. If anyone wants to take this opportunity to work on or to test the trigger system at IR2, please let Selina know.
  • The L3 node trim due to L3 was mentioned in detector operation again last Wednesday. See L3 below for a summary of the node trims.
  • There are TSF damages showing up on June 17 Friday's owl shift (see DCT/DCZ discussion below).
  • We welcomed Fang Fang from Cal-tech joining our Trigger group.

EMT

  • Dan showed us some plots of EMT turn-on for muon-pairs. For the G object, we would need to consider event samples such as radiated bhabha. It would be a few days before Dan commits the codes and puts the plots in OPR or fastmon archive.
  • Rainer and Dan decided to put the EMT prescale factor into database! Dan will look into implementing this. Due to the migration from objectivity to ROOT, it was adviced to delay this a bit later.

DCT/DCZ

  • Two incidences happened to TSF-X15 last Friday. The first one involved the new TSF in run 55381, but the problem only lasted about 10 minutes and then went away even before reconfiguring. The second incident in run 55386 only involved the old TSF system. The new TSF occupancy plots look normal, see plots here, and the dips showed up in the old TSF occupancy plot.) For more details, see HN posted here last Friday.
  • The new online release 13.2.0 will be installed at IR2 before data-taking happens again this week, so we will switch in the new ROM lib that Jamie built against this release. This is happening tomorrow.

L3

Data Quality and Monitoring

  • Kris is reviewing reprocessing of the old data, and one of the runs has a dip in the BLT efficiency (see Page 86 for r12252). According to the elog, this was caused by DCH change of configuration for SL5. Our consent is to mark this run as bad and throw it away.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.